OpenStudy (anonymous):

Write the equation of a parabola with a vertex at (10, 0) and a directrix x = 8. My answer is x-10=1/8y2

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

i think it is not, u must make the graph in order to test ur answer

OpenStudy (anonymous):

I did that was the answer I got

ganeshie8 (ganeshie8):

thats a good guess, you got the focus part wrong. else everything is right

OpenStudy (anonymous):

Is it y-10 then?

OpenStudy (anonymous):

i got y^2 =48(x-10) could u explain the method that u used to see where u made the mistake

OpenStudy (anonymous):

y^2 = 4p(x - h)

ganeshie8 (ganeshie8):

p is the focus ?

OpenStudy (anonymous):

no it isnt typo

ganeshie8 (ganeshie8):

im asking in your equation, y^2 = 4p(x - h), p is the focus, right ?

OpenStudy (anonymous):

O yea p is the focus?

OpenStudy (anonymous):

p is the focus

ganeshie8 (ganeshie8):

ok lets find the focus, then

ganeshie8 (ganeshie8):

your answer, x-10=1/8y2, is correct except for the focus thing. you got h, k correctly there.

ganeshie8 (ganeshie8):